If a cup of ice cream is selected,what is the probability that the cholesterol content will be more than 670 mg?


1
Expert's answer
2022-05-04T17:37:05-0400

Given "\mu=660\ mg, \sigma=35\ mg, n=10"

"P(X>670)=1-P(Z\leq\dfrac{670-660}{35/\sqrt{10}})"




"\approx1-P(Z\leq0.903508)\approx0.183128"

The probability that the mean of the sample will be more than 670 mg is 0.183128.